HUBBLE SPACE TELESCOPE DAILY REPORT #2862 PERIOD COVERED: 0000Z (UTC) 04/30/01 - 0000Z (UTC) 05/01/01 Daily Status Report as of 121/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: The following scheduled proposals did not execute due to the acquisition problems detailed in 2.1 and HSTAR 8196: Three Sets of FGS/1 8731 (A High Angular Resolution Survey of the Most Massive Stars in the SMC) Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) Four Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) Three Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) WF/PC-2 8604 (Stellar Populations Across the Small Magellanic Cloud ( SMC) : History and Structure) STIS/CCD 8865 (Bias Monitor-Part 2) Three Sets of STIS/CCD 8836 (Performance Monitor {Cycle 9}) Six Sets of WF/PC-2 8811 (Cycle 9 Standard Darks) WF/PC-2 8583 (Imaging Snapshots of Asteroids) WF/PC-2 8829 (Observatory Focus Monitor) STIS/CCD 8841 (Hot Pixel Annealing C9) Five Sets of WF/PC-2 8585 (Cosmological Parameters from Type Ia Supernovae at High Redshift) WF/PC-2 8821 (Cycle 9 CTE Monitor) WF/PC-2 8585 (Cosmological Parameters from Type Ia Supernovae at High Redshift) WF/PC-2 8602 (A Snapshot Survey of the Sites of Recent, Nearby Supernovae) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 16 Successful: 8 HSTAR 8196 documents acquisition failures at 120/0208Z, 120/0352Z, 120/0624Z (these three were mentioned in yesterday's report), 120/0907Z, 120/1109Z, 120/1245Z, 120/1421Z, 120/2046Z, 121/0032Z, 121/0215Z, 121/0254Z. Scheduled Re-acquisitions: 1 Successful: 1 2.2 FHST Updates: Scheduled: 22 Successful: 18 HSTAR 8196 mentions update failures at 120/0534Z (covered in yesterday's report), 120/0843Z, 120/1048Z, 120/1228Z, 120/1404Z. 2.3 Operations Notes: Note: Yesterday's report contained an incorrect statement. Part of this section should have read :"Gyro #2 was powered on and gyro #5 was powered OFF at 118/2309Z." Due to the gyro bias update reflecting the new gyro complement that was uplinked at 120/0000Z (HSTAR 8192), there were a series of FGS acquisition and FHST update failures (HSTAR 8196). A series of attitude reference updates (ARUs) were performed but were unsuccessful in regaining FGS control until a new RGA table load was uplinked at 121/0016Z. Another ARU was executed at 121/0343Z and the subsequent FGS acquisitions were successful. Using ROP SR-1A, the SSR EDAC error counter was cleared twice. The CCS engineering status buffer was dumped and cleared at 120/1723Z per ROP DF-18A. A 486 table dump was performed at 120/1454Z as directed by ROP DF-01A. A TTR was written when an unexpected TDE support at 120/2247Z became available. Another TTR detailed the TDE support at 120/2315Z for which 171 support could not be scheduled. As a result there was a 25-minute loss of engineering data. Using ROP DF-07A, the ephemeris load was uplinked at 121/0350Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW